OpenStudy (anonymous):

How do I find the Y interept to the following function? P(x) = (x − 5)^3(x − 2)^2(x +1)

OpenStudy (anonymous):

intercept

OpenStudy (aravindg):

Put x=0 in the equation to get y intercept.

OpenStudy (anonymous):

(0-5)^3(0-2)^2(0+1)

OpenStudy (anonymous):

P(x)= -125+4+1 =

OpenStudy (anonymous):

did i calculate that right? I don't think I did

OpenStudy (anonymous):

-5(-2)(1) = 10 or is that ?

OpenStudy (aravindg):

It will be (-5)^3 times (-2)^2

OpenStudy (aravindg):

You forgot the exponents

OpenStudy (anonymous):

P(x)= -125+4+1 ?

OpenStudy (anonymous):

i don't think that's right

OpenStudy (aravindg):

(-125)(4)(1)

OpenStudy (anonymous):

ohhhh okay lol that makes sense thank you

OpenStudy (aravindg):

LOL FINALLY :D